# Apollo: An Lightweight Multilingual Medical LLM towards Democratizing Medical AI to 6B People

### v2

@@ Appendix B Models
Qwen is a suite of large language models from the Aliyun-developed Tongyi Qianwen from 0.5 billion to 72 billion parameters, based on the Transformer architecture and are trained on a diverse and extensive range of pretraining data. The types of pretraining data are varied and cover a wide scope, including a vast array of internet texts, professional books, code, and more.
Meditron
@@
Meditron is a suite of open-source medical large language models from 7 billion to 70 billion parameters, adapted to the medical domain from Llama-2 through continued pretraining on a comprehensively curated medical corpus, including selected PubMed articles, abstracts, a new dataset of internationally-recognized medical guidelines, and general domain data from RedPajama-v1.
Llama-2
@@
Llama-2 is a collection of pretrained and fine-tuned generative text models ranging in scale from 7 billion to 70 billion parameters. The fine-tuned versions use supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align to human preferences for helpfulness and safety.
Gemma
@@
Gemma is a family of lightweight, state-of-the-art open models from Google, built from the same research and technology used to create the Gemini models. They are text-to-text, decoder-only large language models, available in English, with open weights, pre-trained variants, and instruction-tuned variants.
MMedLM2
@@
MMedLM 2 is a multilingual medical foundation model available in two versions, with parameter sizes of 1.8 billion and 7 billion. MMedLM 2 builds upon the foundation of InternLM 2 and has been further pretrained on MMedC, a comprehensive multilingual medical corpus. This further pretraining enhances the model’s medical-domain knowledge.
Yi
@@
The Yi series models are the next generation of open-source large language models trained from scratch by 01.AI. Targeted as a bilingual language model and trained on 3T multilingual corpus, the Yi series models show promise in language understanding, commonsense reasoning, reading comprehension, and more.
Mistral
@@
Mistral is a pretrained generative text model with 7 billion parameters. It uses Grouped-query attention (GQA) for faster inference and Sliding Window Attention (SWA) to handle longer sequences at smaller cost.
Zephyr
@@
Zephyr is a series of language models that are trained to act as helpful assistants, which is a fine-tuned version of mistralai/Mistral-7B-v0.1 that was trained on on a mix of publicly available, synthetic datasets using Direct Preference Optimization (DPO).
BioMistral
@@
BioMistral is a suite of Mistral-based further pre-trained open source models suited for the medical domains and pre-trained using textual data from PubMed Central Open Access. All the models are trained using the CNRS (French National Centre for Scientific Research) Jean Zay French HPC.
HuatuoGPT-2
@@
HuatuoGPT2 is a suite of open-source medical large language models from 7 billion to 34 billion parameters, which employs an innovative domain adaptation method to significantly boost its medical knowledge and dialogue proficiency. It showcases state-of-the-art performance in several medical benchmarks, especially surpassing GPT-4 in expert evaluations and the fresh medical licensing exams.
PMC-Llama
@@
MedLlama is initialized from Llama-13B and further pretrained with medical corpus. Despite the expert knowledge gained, it lacks instruction-following ability. It provides a instruction-tuning dataset and evaluates the tuned model. MedLlama is pretrained on medical corpus, and PMC_Llama is further finetuned based on MedLlama.
Apollo (Ours)
@@
Apollo is a suite of open-source medical large language models from 1.8 billion to 7 billion parameters. The priority of all data items from the pre-training corpus to 16, and the priority of all data items from the Instruction Tuning stage to 2. The Batch size of model training is set to 256, the learning rate is set to 1e-4 for most models and 1e-5 for 7B model, and the warm up rate of Cosine scheduler is set to 0.03. The pre-training corpus is trained for one epoch, the Instrument Tuning corpus is trained for two epochs.

### v4

@@ B.2 Models for XMedBench
Qwen Qwen is a suite of large language models from the Aliyun-developed Tongyi Qianwen from 0.5 billion to 72 billion parameters, based on the Transformer architecture and are trained on a diverse and extensive range of pretraining data. The types of pretraining data are varied and cover a wide scope, including a vast array of internet texts, professional books, code, and more.
@@
Meditron Meditron is a suite of open-source medical large language models from 7 billion to 70 billion parameters, adapted to the medical domain from Llama-2 through continued pretraining on a comprehensively curated medical corpus, including selected PubMed articles, abstracts, a new dataset of internationally-recognized medical guidelines, and general domain data from RedPajama-v1.
@@
Llama-2 Llama-2 is a collection of pretrained and fine-tuned generative text models ranging in scale from 7 billion to 70 billion parameters. The fine-tuned versions use supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align to human preferences for helpfulness and safety.
@@
Gemma Gemma is a family of lightweight, state-of-the-art open models from Google, built from the same research and technology used to create the Gemini models. They are text-to-text, decoder-only large language models, available in English, with open weights, pre-trained variants, and instruction-tuned variants.
@@
MMedLM2 MMedLM 2 is a multilingual medical foundation model available in two versions, with parameter sizes of 1.8 billion and 7 billion. MMedLM 2 builds upon the foundation of InternLM 2 and has been further pretrained on MMedC, a comprehensive multilingual medical corpus. This further pretraining enhances the model’s medical-domain knowledge.
@@
Yi The Yi series models are the next generation of open-source large language models trained from scratch by 01.AI. Targeted as a bilingual language model and trained on 3T multilingual corpus, the Yi series models show promise in language understanding, commonsense reasoning, reading comprehension, and more.
@@
Mistral Mistral is a pretrained generative text model with 7 billion parameters. It uses Grouped-query attention (GQA) for faster inference and Sliding Window Attention (SWA) to handle longer sequences at smaller cost.
@@
Zephyr Zephyr is a series of language models that are trained to act as helpful assistants, which is a fine-tuned version of mistralai/Mistral-7B-v0.1 that was trained on on a mix of publicly available, synthetic datasets using Direct Preference Optimization (DPO).
@@
BioMistral BioMistral is a suite of Mistral-based further pre-trained open source models suited for the medical domains and pre-trained using textual data from PubMed Central Open Access. All the models are trained using the CNRS (French National Centre for Scientific Research) Jean Zay French HPC.
@@
HuatuoGPT-2 HuatuoGPT2 is a suite of open-source medical large language models from 7 billion to 34 billion parameters, which employs an innovative domain adaptation method to significantly boost its medical knowledge and dialogue proficiency. It showcases state-of-the-art performance in several medical benchmarks, especially surpassing GPT-4 in expert evaluations and the fresh medical licensing exams.
@@
PMC-Llama MedLlama is initialized from Llama-13B and further pretrained with medical corpus. Despite the expert knowledge gained, it lacks instruction-following ability. It provides a instruction-tuning dataset and evaluates the tuned model. MedLlama is pretrained on medical corpus, and PMC_Llama is further finetuned based on MedLlama.
@@
Apollo (Ours) Apollo is a suite of open-source medical large language models from 1.8 billion to 7 billion parameters. The priority of all data items from the pre-training corpus to 16, and the priority of all data items from the instruction tuning stage to 2. The Batch size of model training is set to 256, the learning rate is set to 1e-4 for most models and 1e-5 for 7B model, and the warm up rate of Cosine scheduler is set to 0.03. The pre-training corpus is trained for one epoch, the Instrument Tuning corpus is trained for two epochs.
Appendix C Settings of Proxy-Tuning